OpenStudy (anonymous):

when you're subtracting rationals, do you make it into a plus sign and make EVERYTHING in the numerator of the second negative, or just make 1 number negative? for examples x-1 4x+2 ---- - ----- x(x-3) x(x-3) would you make it into a plus sign, and then make 4x and 2 negative? -4x-2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

in the other equation you did how did you get positive 2?

OpenStudy (anonymous):

the last one you answered

OpenStudy (anonymous):

\[\frac{x-1-(4x+2)}{x(x-3)}\] \[\frac{x-1-4x-2}{x(x-3)}\] etc

OpenStudy (anonymous):

well in the last one you caught me in a dumb algebra mistake that is why!

OpenStudy (anonymous):

should be \[\frac{x+2}{x}=\frac{x-1}{x}-\frac{4x+2}{x(x-3)}\] \[(x+2)(x-3)=(x-1)(x+3)-4x-2\] \[x^2-x-6 = x^2-2 x-5\] \[-x-6=-2x-5\] \[x=1\]

OpenStudy (anonymous):

wait what? i get x^2-x-6=x^2-8x+1 i keep on looking it over but i can't figure it out darn it

OpenStudy (anonymous):

oh man did i mess up again? hold on

OpenStudy (anonymous):

ok so, (x-3) and (x+2) equal x^2-x-6. (x-1) and (x-3) euqlas x^2-4x+3. then when i combine them x^2-x-6=x^2-4x+3x-4x-2

OpenStudy (anonymous):

ok now you have caught me in two mistakes, so i am going to quit. you are right it is \[(x+2)(x-3)=(x-1)(x-3)-4x-2\]

OpenStudy (anonymous):

then i combine like terms and get x^2-x-6=x^2-8x+1

OpenStudy (anonymous):

no please don't this is a very confusing problem lol

OpenStudy (anonymous):

then i get 2x^1+7x+7

OpenStudy (anonymous):

ok but i am making mistake after mistake. lets see what we get now

OpenStudy (anonymous):

i mean 2x^2+7x+7

OpenStudy (anonymous):

did you get \[x^2-x-6 = x^2-8 x+1\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok subtract \[x^2\] from both sides

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

7x+7

OpenStudy (anonymous):

what happened to your equal sign?

OpenStudy (anonymous):

0=7x+7

OpenStudy (anonymous):

oh wait let me do this again

OpenStudy (anonymous):

OMFG

OpenStudy (anonymous):

\[x^2-x-6 = x^2-8 x+1\] \[\cancel{x^2}-x-6=\cancel{x^2}-8x+1\] \[-x-6=-8x+1\]

OpenStudy (anonymous):

I GOT THE RIGHT ANSWER THANK YOU!!!!!!!!!!!!!! FINALLY THIS PROBLEM IS THE WORST THANK YOU!!!!!!!!!!!!!

OpenStudy (anonymous):

YES!!!!!!!!!

OpenStudy (anonymous):

took forever!!!!!!!

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

yw. now i have a question. what is vivistee?

OpenStudy (anonymous):

my cat that ran away

OpenStudy (anonymous):

never forget

OpenStudy (anonymous):

listen thanks alright? my gosh that problem was so confusing and frustrating

OpenStudy (anonymous):

:"}

OpenStudy (anonymous):

it's finally over now :"}

OpenStudy (anonymous):

i'm going to write that one down in my notes

OpenStudy (anonymous):

glad you got it, sorry about your cat